# Ensemble Kalman Filtering-Aided Variational Inference for Gaussian Process State-Space Models

### 10 Dec 2023 v1

@@ Abstract
Gaussian process state-space models (GPSSMs) provide a principled and flexible approach to model latent state dynamics observed through emission models. However, existing variational methods for learning GPSSMs face a substantial challenge in optimizing a large number of parameters, particularly with the introduction of amortized inference networks. To address this challenge, we leverage the ensemble Kalman filter (EnKF), a well-established model-based filtering technique, to approximate the posterior distribution of latent states within the variational inference framework. This approach eliminates the need for inference networks, significantly reducing the number of variational parameters. Moreover, we demonstrate that with the aid of EnKF, the straightforward evaluation of approximated evidence lower bound (ELBO) in the variational inference can be easily obtained through the summation of multiple terms with closed-form solutions. By leveraging automatic differentiation tools, we thus can maximize the ELBO and train the GPSSM efficiently. We also extend the proposed method to an online setting and provide comprehensive algorithm analyses and insights. Extensive testing on diverse real and simulated datasets demonstrates that our variational inference algorithms, integrated with EnKF, outperform existing methods in terms of learning and inference performance.
@@
Gaussian process, state-space model, ensemble Kalman filter, online learning, variational learning and inference.
@@ I Introduction
State-space models (SSMs) are versatile modeling tools used to characterize dynamical patterns in time series data [1]. They have demonstrated successful applications in a wide range of fields, including engineering, statistics, computer science, and economics. The primary tasks involved in SSMs are learning and inference [2]. Learning, also known as system identification, focuses on determining optimal model parameters to accurately represent the underlying dynamical system. In contrast, inference aims to estimate latent states of interest based on observed sequential data [1, 2, 3].
@@
Over the past few decades, numerous learning and inference methods have been developed for SSMs. In cases where the system dynamics are precisely known, various well-established techniques such as the Kalman filter (KF), extended Kalman filter (EKF), ensemble Kalman filter (EnKF) and particle filter (PF) can be employed to infer the latent states [1]. However, in complex and uncertain scenarios like model-based reinforcement learning [4, 5] and disease epidemic propagation [6], knowing and determining the underlying system dynamics in advance poses tremendous challenges [7]. Thus, the dynamics need to be learned from the observed noisy measurements, leading to the emergence of data-driven state-space models. One class of prominent data-driven state-space models is the Gaussian process state-space model (GPSSM) [8], which utilizes Gaussian processes (GPs) as core components to capture the complex dynamics of the underlying system [9].
The popularity of GPSSM is largely driven by its inherent appealing properties [8]. First, GPSSM is a probabilistic SSM, offering interpretability and explicit uncertainty calibration for the system dynamics [10]. Second, the nonparametric GP models in GPSSM enable automatic adaptation to problem complexity. Consequently, in practice, GPSSMs often achieve high accuracy in model learning even with limited data [2].
Despite the numerous advantageous properties of GPSSMs, the task of learning and inference in GPSSMs remains challenging. Early studies primarily focused on the following two approaches: (1) inferring the latent states through maximum a posteriori (MAP) estimation [10, 11, 12, 13], and/or (2) learning GPSSM transition function by assuming the observable latent states [14, 15, 16]. However, the MAP estimation is limited to providing a point estimate and often fails to accurately capture the mode of the posterior distribution [3], while assuming observable latent states for learning transition function is often impractical in various applications [17]. These challenges have led to a limited applicability of the earlier GPSSM works.
@@
To overcome these challenges, the seminal work in [18] introduces a fully Bayesian framework, utilizing Monte Carlo sampling techniques for simultaneous learning and inference of the GPSSM. Subsequently, more advanced approximation methods [19, 20, 21] leverage the GP approximation from [22] to alleviate the substantial computational demands in [18]. However, the computational burden associated with Monte Carlo sampling methods, particularly when dealing with high-dimensional latent states, remains notably significant. Consequently, there has been a shift towards variational inference methods [23, 24, 25, 26, 27, 28, 29] that combine with the utilization of sparse GPs featuring inducing points [30]. Typically, the variational inference approaches can be classified into two categories: the mean-field (MF) class [23, 8, 24, 25], and the non-mean-field (NMF) class [26, 27, 28, 29], based on whether the independence assumption is applied to the variational distribution concerning the GP-based transition function and latent state.
The first MF variational algorithm that incorporates a PF for latent state inference is introduced in [23]. To address the computational complexity inherent in the PF, the work in [24] proposes the utilization of a (parametric) linear Markov Gaussian variational distribution for the latent states. However, this approach presents an issue of linear growth in the number of variational parameters over time. In response, the authors in [25] present a solution by leveraging an amortized inference network [31] to model the variational distribution effectively. In subsequent research endeavors, the focus has shifted towards the NMF algorithm, aiming to improve inference accuracy and mitigate issues of over-confidence [32].
The first NMF variational algorithm is proposed in [26], where the posterior distribution of the latent states conditioning on the GP transition function is directly approximated using the prior distribution. However, this approximation lacks the incorporation of observation information in the posterior, rendering filtering or smoothing infeasible. To address this issue, the work in [27] explicitly assumes a flexible parametric Markov-structured Gaussian variational distribution over the latent states, and propose an algorithm called VCDT (variationally coupled dynamics and trajectories). In [28], the authors propose a forward-backward smoothing-like variational learning method to address learning in unstable systems with partial observability. In [29], the Laplace approximation is employed to approximate the posterior distribution, resulting in a linear Markov Gaussian approximation similar to [25]. Other recent works have highlighted the advantageous impact of increased GPSSM’s flexibility on learning and inference performance [33, 34, 35, 36, 37].
Nevertheless, the existing methods suffer from two significant limitations. First, the state-of-the-art NMF assumption on the variational distribution could still be oversimplified, which may overlook the intricacies of the underlying dynamics [38]. Second, optimizing the amortized inference network in state-of-the-art methods [25, 26, 27, 28, 36] presents notable computational challenges due to the substantial data requirements and the inherent risk of encountering numerous bad local optima during the optimization process [2]. In response to the first limitation, a very recent work [38] proposes using a free-form variational distribution in combination with stochastic gradient Hamiltonian Monte Carlo [39]. In this paper, we focus on the second limitation. Specifically, our objective is to alleviate the optimization and computational burden by eliminating the need for additional variational parameters and black-box inference networks. The main contributions of this paper are summarized as follows.
@@
  • We highlight the significance of incorporating well-established model-based filtering techniques in developing GPSSM variational algorithms. In contrast to existing methods relying on inference networks with numerous variational parameters for learning variational distribution, we introduce a novel EnKF-aided NMF variational inference algorithm tailored for GPSSMs. This novel approach eliminates the need for an inference network, resulting in a substantial reduction in variational parameters, thereby enhancing learning efficiency.
@@
  • In addition, with the aid of EnKF, we demonstrate that the straightforward evaluation of approximated evidence lower bound (ELBO) in the variational inference can be easily obtained by summing multiple terms with closed-form analytical solutions. By harnessing automatic differentiation tools, we thus can efficiently optimize the ELBO and train the GPSSM.
@@
  • Without the burden of inference network, the proposed EnKF-aided algorithm gets readily extended to an online learning setting. This online extension introduces a more comprehensive objective function, making it superior to state-of-the-art online algorithms in learning and inference efficiency.
@@
  • We provide comprehensive analysis and profound insights into the proposed algorithms and conduct extensive experiments on various real and synthetic datasets to evaluate the algorithms’ performance. The results demonstrate that the proposed EnKF-aided variational learning methods consistently outperform existing state-of-the-art methods in terms of learning and inference performance.
@@ III Problem Statement
See detailed derivations in Supplement D-A. Maximizing $\mathcal{L}$ w.r.t. $\bm{\theta}$ fine-tunes the model parameters to improve the fit with the observed data, reflecting the model learning process; while maximizing $\mathcal{L}$ w.r.t. the variational distribution $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$ is equivalent to minimizing the KL divergence, $\operatorname{KL}[q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})\|p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})]$. That is, it enhances the quality of the variational distribution approximation, bringing it closer to the underlying posterior distribution, $p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})$, which corresponds to the model inference process [40]. The capacity of $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$ to approximate $p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})$ is thus of crucial importance [41].
@@
where $q({\vec{\mathbf{f}}})$ represents the variational distribution of the GP and $q({\mathbf{x}}_{0})\prod_{t=1}^{T}q({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$ corresponds to the variational distribution of the latent states [42, 36]. It is noteworthy that the factorization of the variational distribution presented in Eq. (8) is recognized as a NMF approximation within the GPSSM literature (see formal definition in Appendix C) [27], because it explicitly establishes the dependence between the latent states and the GP transition function values, as manifested by the terms $\prod_{t=1}^{T}q({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$.
The first notable NMF approximation approach is proposed by Doerr et al. in [26], where $q({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$ is set equal to the prior distribution $p({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$, i.e., $q({\mathbf{x}}_{t}|{\mathbf{f}}_{t})=p({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$. Although there is a dependence between latent states ${\vec{\mathbf{x}}}$ and transition function values ${\vec{\mathbf{f}}}$, there is no filtering or smoothing effect for latent states [27]. Instead, Ialongo et al. assume a flexible, parametric Markov-structured Gaussian variational distribution over the temporal states, i.e.,
@@
where $\{\bm{A}_{t},\bm{b}_{t},\bm{S}_{t}\}_{t=1}^{T}$ are free variational parameters. This choice, however, introduces a significant drawback: the number of variational parameters grows linearly with the length of the time series. Although this issue can be partially addressed by incorporating an inference network, such as a bidirectional recurrent neural network (Bi-RNN) [25, 42], to amortize the inference process, fine-tuning the large variational parameters in the inference network still requires substantial effort. Consequently, despite the flexibility of the variational distribution described in Eq. (9) allowing for approximation of the true posterior distribution, its empirical performance often falls short of its theoretical expressive capacity [29]. Moreover, it is worth noting that the inference network commonly takes the input of the observations, ${\vec{\mathbf{y}}}$, rendering the existing inference network-based methods predominantly trained in an offline manner, thus posing challenges when attempting to adapt them for online applications. To address these limitations, this paper proposes a novel approach that combines well-established EnKF with modern automatic differentiation tools [43] for variational inference in GPSSMs. The objective is to alleviate the optimization and computational burden, as well as to develop a more interpretable approach, stepping away from black-box inference networks.
@@ IV EnVI: EnKF-Aided Variational Inference
This section presents our novel variational inference algorithm for GPSSMs. We begin by introducing the model-based EnKF in Section IV-A, which serves as the foundation for our algorithm. Following that, Section IV-B details the proposed EnKF-aided variational inference algorithm. Lastly, extensive discussions about the properties of our proposed method are given in Section IV-C.
@@ IV-A Ensemble Kalman Filter (EnKF)
Given a SSM, see Eq. (II-B), the EnKF sequentially approximates the filtering distributions using $N\in\mathbb{N}$ equally weighted particles [44]. Specifically, at the prediction step, EnKF first samples particles, ${\mathbf{x}}_{t-1}^{1:N}\triangleq\{{\mathbf{x}}_{t-1}^{n}\}_{n=1}^{N}$, from the filtering distribution at time $t-1$, $p({\mathbf{x}}_{t-1}|{\mathbf{y}}_{1:t-1})$. Then, the sampled particles are propagated using the state transition function $f(\cdot)$, i.e.,
@@
With the linear Gaussian emission in Eq. (II-B), the joint distribution of ${\mathbf{x}}_{t}$ and ${\mathbf{y}}_{t}$ can be expressed as follows:
@@
Thus, at the update step, the filtering distribution, $p({\mathbf{x}}_{t}|{\mathbf{y}}_{1:t})={\cal N}({\mathbf{x}}_{t}|\mathbf{m}_{t},\mathbf{P}_{t}),$ can be obtained using conditional Gaussian identity, where
@@
In the following subsections, we will describe our novel integration of the variational inference framework with the EnKF [44], which leads to a reduction in variational parameters and ultimately enhances the learning and inference performance of GPSSMs.

### 23 Dec 2023 v2

@@ Abstract
Gaussian process state-space models (GPSSMs) are a flexible and principled approach for modeling dynamical systems. However, existing variational learning and inference methods for GPSSMs often necessitate optimizing a substantial number of variational distribution parameters, leading to inadequate performance and efficiency. To overcome this issue, we propose incorporating the ensemble Kalman filter (EnKF), a well-established model-based filtering technique, into the variational inference framework to approximate the posterior distribution of latent states. This utilization of EnKF can effectively exploit the dependencies between latent states and GP dynamics, while eliminating the need for parameterizing the variational distribution, thereby significantly reducing the number of variational parameters. Moreover, we show that our proposed algorithm allows straightforward evaluation of an approximated evidence lower bound (ELBO) in variational inference via simply summating multiple terms with readily available closed-form solutions. Leveraging automatic differentiation tools, we hence can maximize the ELBO and train the GPSSM efficiently. We also extend the proposed algorithm to an online setting and provide detailed algorithmic analyses and insights. Extensive evaluation on diverse real and synthetic datasets demonstrates the superiority of our EnKF-aided variational inference algorithms in terms of learning and inference performance compared to existing methods.
@@
Gaussian process, state-space model, ensemble Kalman filter, online learning, variational inference.
@@ I Introduction
State-space models (SSMs) describe the underlying dynamics of latent states through a transition function and an emission function [1]. As a versatile tool for modeling dynamical systems, SSM finds successful applications in diverse fields, including control engineering, signal processing, computer science, and economics [2, 3, 4]. In SSM, a key task involves inferring unobserved latent states from a sequence of noisy measurements. Established techniques, such as the Kalman filter (KF), extended Kalman filter (EKF), ensemble Kalman filter (EnKF), and particle filter (PF), have been widely employed over the decades for latent state inference [2]. However, these state inference methods heavily rely on precise knowledge of the underlying system dynamics [5], posing tremendous challenges in complex and uncertain scenarios like model-based reinforcement learning [6, 7] and disease epidemic propagation [8]. Consequently, the underlying dynamics need to be learned from noisy observations, leading to the emergence of data-driven state-space models. One class of prominent data-driven state-space models is the Gaussian process state-space model (GPSSM) [9], which utilizes Gaussian processes (GPs) as core learning modules to capture the complex underlying system dynamics [10].
Gaussian processes, serving as a Bayesian nonparametric model [11], provide the flexibility to model general nonlinear system dynamics without enforcing an explicit parametric structure. With the inherent regularization imposed by the GP prior, GPSSMs are able to mitigate model overfitting [12], rendering them more effective in scenarios with limited data [13]. Moreover, inheriting the Bayesian nature from GPs, GPSSMs offer interpretability and explicit uncertainty calibration for the system dynamics. As a result, GPSSMs exhibit increased transparency and caution in regions of the state space where data is scarce, which is crucial for safety-critical applications [14]. Due to these advantageous properties, GPSSMs have seen extensive use in various applications, such as human pose and motion learning [14], robotics and control learning [15, 16, 17], reinforcement learning [18], target tracking and navigation [19], and magnetic-field estimation [20].
Despite the popularity of GPSSMs, simultaneously learning the model and estimating the latent states in GPSSMs remains highly challenging, primarily due to the following two factors. First, the quality of state estimates affects model learning and vice versa, which introduces computational and statistical complexities. Second, the nonlinearity in GPSSMs renders the learning and inference analytically intractable [11]. Hence, the main goal in GPSSM is to accurately approximate the joint posterior over latent states and GP-modeled system dynamics. Multiple approaches have emerged towards achieving this objective in the last decade.
@@
The seminal work utilizing particle Markov chain Monte Carlo (PMCMC) is proposed in [9]. Subsequently, more advanced methods [21, 22, 23] leverage the reduced-rank GP approximation introduced in [24] to alleviate the substantial computational demands of GPs in [9]. However, the computational burden of the involved PMCMC remains notably significant, particularly when dealing with high-dimensional latent states. Consequently, there has been a shift towards variational inference methods [25, 26, 27, 28, 29, 30, 31, 32, 33], which utilize the sparse GP approximations with inducing points [34].
Generally, the variational inference approaches can be classified into two categories: the mean-field (MF) class [25, 26, 27, 28], and the non-mean-field (NMF) class [29, 30, 31, 32, 33], based on whether the independence assumption is applied to the variational distribution concerning latent states and GP dynamics. The first variational algorithm, incorporating the MF assumption and integrating PF for latent state inference, is introduced in [25]. Subsequent MF-based endeavors [26, 27, 28] aim at reducing computational and model complexities in [25]. Yet, the MF class method neglects the dependencies between latent states and GP dynamics, potentially compromising learning accuracy and yielding overconfident state estimates [35].
More recent methods concentrate on NMF variational approximations [29, 30, 31, 32, 33]. Specifically, in [29], the posterior distribution of underlying states conditioned on the GP dynamics is directly approximated by the corresponding prior distribution; while in [30] and [31], the posterior is approximated by parametric Markov-structured Gaussian variational distributions. In contrast to the heavy parameterization as in [30] and [31], the works in [32] and [33] opt for directly characterizing the posterior distribution through the model evidence. However, the Laplace approximation in [32] can oversimplify the approximated distribution over latent states (unimodal approximation) [11], and the stochastic gradient Hamiltonian Monte Carlo method [36] employed in [33] exhibits slow convergence ($50,000$ iterations as reported). Other recent works, such as [37, 38, 39, 40], have highlighted the advantageous impact of increased GPSSM’s flexibility on learning and inference performance.
Nevertheless, the existing methods face two significant issues when approximating the joint posterior distribution concerning latent states and GP dynamics. First, the majority of existing approaches [27, 28, 29, 30, 31, 33, 38, 39, 40] continue to struggle with computational challenges, manifested by the extensive optimization parameters and notably slow convergence of the optimization algorithms. Additionally, the existing variational methods are incompatible with online applications, posing a notable limitation in their applicability.
To address the two issues mentioned above, our proposed methods focus on two critical aspects in terms of the joint posterior distribution approximation: 1) whether the dependencies between GP dynamics and latent states are fully exploited during algorithm design and 2) whether the corresponding variational distributions are as less parameterized as possible. We show that by exploiting the dependencies between latent states and GP dynamics via EnKF, our methods can significantly alleviate the model and computational complexity and get readily extended to the online setting. The main contributions of this paper are summarized as follows.
@@
  • We highlight the significance of incorporating well-established model-based filtering techniques to enhance the efficiency of GPSSM variational algorithms. In contrast to existing methods utilizing numerous variational parameters to parameterize the variational distribution over latent states, our proposed novel EnKF-aided NMF variational inference algorithm harnesses the EnKF to exploit the dependencies between latent states and GP dynamics, and eliminate the need to parameterize the variational distribution, significantly reducing the number of variational parameters and accelerating the algorithm convergence.
@@
  • By leveraging EnKF, we also demonstrate the simplified computation of the approximated evidence lower bound (ELBO) in variational inference, by aggregating multiple terms with readily available closed-form analytical solutions. Leveraging the differentiability inherent in EnKF alongside automatic differentiation tools, we can efficiently optimize the ELBO and train the GPSSM.
@@
  • Without explicitly parameterizing the variational distribution over latent states, the proposed EnKF-aided algorithm gets readily extended to an online learning setting. This online extension introduces a more comprehensive objective function, making it superior to state-of-the-art online algorithms in learning and inference efficiency.
@@
  • We provide detailed analysis and fresh insights of the proposed algorithms and conduct extensive experiments on various real and synthetic datasets to evaluate the algorithms’ performance. The results demonstrate that the proposed EnKF-aided variational learning methods consistently outperform existing state-of-the-art methods in terms of learning and inference performance.
@@ III Problem Statement
Detailed derivations can be found in Supplement D-A. Maximizing $\mathcal{L}$ w.r.t. $\bm{\theta}$ fine-tunes the model parameters to improve the fit with the observed data, reflecting the model learning process; while maximizing $\mathcal{L}$ w.r.t. the variational distribution $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$ is equivalent to minimizing the KL divergence, $\operatorname{KL}[q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})\|p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})]$. That is, it enhances the quality of the variational distribution approximation, bringing it closer to the underlying posterior distribution, $p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})$, which corresponds to the model inference process [41]. The capacity of $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$ to approximate $p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})$ is thus of crucial importance [42].
@@
where $q({\vec{\mathbf{f}}})$ represents the variational distribution of the GP and $q({\mathbf{x}}_{0})\prod_{t=1}^{T}q({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$ corresponds to the variational distribution of the latent states [43, 39]. It is noteworthy that the factorization of the variational distribution presented in Eq. (8) is recognized as a NMF approximation within the GPSSM literature (see formal definition in Appendix C) [30], because it explicitly establishes the dependencies between the latent states and the GP transition function values, as manifested by the terms $\prod_{t=1}^{T}q({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$.
The methods closely related to this paper are probably [29] and [30]. In [29], variational distribution $q({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$ is set equal to the prior distribution $p({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$, i.e., $q({\mathbf{x}}_{t}|{\mathbf{f}}_{t})=p({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$. But this choice results in no filtering or smoothing effect on latent states, although there is a dependence between latent states ${\vec{\mathbf{x}}}$ and transition function values ${\vec{\mathbf{f}}}$. Instead, the work in [30] assumes a parametric Markov-structured Gaussian variational distribution over the temporal states, i.e.,
@@
where $\{\bm{A}_{t},\bm{b}_{t},\bm{S}_{t}\}_{t=1}^{T}$ are free variational parameters. This choice, however, introduces a significant drawback: the number of variational parameters grows linearly with the length of the time series. Although this issue can be partially addressed by incorporating an inference network, such as a bidirectional recurrent neural network [28, 43], to learn the variational parameters and make the variational distribution learning with a constant number of parameters, fine-tuning these variational parameters in the inference network still requires substantial effort. Consequently, despite the flexibility of the variational distribution described in Eq. (9) allowing for approximation of the true posterior distribution, its empirical performance often falls short of its theoretical expressive capacity [32]. Moreover, it is worth noting that the inference network commonly takes the input of the observations, ${\vec{\mathbf{y}}}$, rendering the existing inference network-based methods predominantly trained in an offline manner, thus posing challenges when attempting to adapt them for online applications. To address these limitations, this paper proposes a novel approach that combines well-established EnKF with modern automatic differentiation tools [44] for variational inference in GPSSMs. The main objective is to alleviate the model and computational complexity and develop a more interpretable approach (e.g., stepping away from the potential black-box inference networks).
@@ IV EnVI: EnKF-Aided Variational Inference
This section presents our novel variational inference algorithm for GPSSMs. We begin by introducing the model-based EnKF in Section IV-A, which serves as the foundation for our algorithm. Following that, Section IV-B details the proposed EnKF-aided variational inference algorithm. Lastly, extensive discussions about the properties of our proposed algorithm are given in Section IV-C.
@@ IV-A Ensemble Kalman Filter (EnKF)
Given a SSM, see Eq. (II-B), the EnKF sequentially approximates the filtering distributions using $N\in\mathbb{N}$ equally weighted particles [45]. Specifically, at the prediction step, EnKF first samples particles, ${\mathbf{x}}_{t-1}^{1:N}\triangleq\{{\mathbf{x}}_{t-1}^{n}\}_{n=1}^{N}$, from the filtering distribution, $p({\mathbf{x}}_{t-1}|{\mathbf{y}}_{1:t-1})$, at time $t-1$. Then, the sampled particles are propagated using the state transition function $f(\cdot)$, i.e.,
@@
With the linear Gaussian emission in Eq. (II-B), the joint distribution of ${\mathbf{x}}_{t}$ and ${\mathbf{y}}_{t}$ can be readily obtained as follows:
@@
Thus, the filtering distribution, $p({\mathbf{x}}_{t}|{\mathbf{y}}_{1:t})\!=\!{\cal N}({\mathbf{x}}_{t}|\mathbf{m}_{t},\mathbf{P}_{t}),$ can be obtained using conditional Gaussian identity at the update step, where
@@
In the following subsections, we will describe our novel integration of the variational inference framework with the EnKF [45], which leads to a reduction in the number of variational parameters and ultimately enhances the learning and inference performance of GPSSMs.